-2x -8y = -3
-10x -2y = -6
This is from a CNOW project.
If the system is to be solved using the addition method, by which constant should one of the equations be multiplied if the x terms are to drop?
-5 , 2 , 1/5 , or -2

Answer by stanbon(75887)

-2x -8y = -3
-10x -2y = -6
This is from a CNOW project.
If the system is to be solved using the addition method, by which constant should one of the equations be multiplied if the x terms are to drop?
-5 , 2 , 1/5 , or -2
-------------
Ans: -5
============
Cheers,
Stan H.